OpenStudy (anonymous):

Suppose you wish to buy a DVD that cost $9 and the sales tax rate is 6.5%. Calculate the amount of sales tax and the total price you will pay for the DVD. Select one: a. $0.58, $9.58 Incorrect b. $0.59, $9.58 c. $0.59, $9.59 d. $0.60, $9.60

OpenStudy (anonymous):

@TSwizzle

OpenStudy (campbell_st):

tax = 9 x 6.5/100 price = 9 + tax

OpenStudy (anonymous):

Is That An Add Sign Or Division

OpenStudy (campbell_st):

\[tax = 9 \times 6.5 \div 100\] \[price = 9 + tax\]

OpenStudy (anonymous):

It Should Be 9.585 But Its None Of The Options